Consider a four-step serial process with processing times given in the following list. There is one machine at each step of the process, and this is a machine-paced process. bullet
Step 1: 25 minutes per unit bullet
Step 2: 15 minutes per unit bullet
Step 3: 30 minutes per unit bullet
Step 4: 20 minutes per unit
Assuming that the process starts out empty, how long will it take (in hours) to complete a batch of 105 units?
Answer #1

Solution :

The rate of system is defined by machine with largest time per unit i.e. bottleneck resource.

In this case, biottleneck is Step : 3 with time 30 min per unit.

So system requires 30 * 105 min = 30 * 105 / 60 hours = 52.5 hours at bottleneck.

time for first unit at Step 1 and 2 and for last unit at step 4 is - 25 + 15 + 20 - 60min - 1 hr.

So total time for system is 52.5 + 1 - 53.5 hours.